OpenStudy (chris215):

Find f(g(-3)) if 2/(x+4) and g(x) = (x - 1)^2

OpenStudy (chris215):

How do I do this?

OpenStudy (anonymous):

First find g(-3) by plugging x = -3 into g(x)

OpenStudy (chris215):

I plugged x=-3 into g(x) = (x - 1)^2 and got 16

OpenStudy (chris215):

Find f(g(-3)) if f(x)= 2/(x+4) and g(x) = (x - 1)^2 im sorry I forgot the f(x)

OpenStudy (anonymous):

Awesome Now, f(g(-3) = f(16) Can you find f(16) ?

OpenStudy (chris215):

1/10

OpenStudy (chris215):

Thanks so much !

OpenStudy (anonymous):

You are welcome :)